# Furstenberg set problem and exceptional set estimate in prime fields: dimension two implies higher dimensions

### v1

@@ 5. Furstenberg set problem: Lower bound II
Return back to $E_{W}:=\bigcup_{V\in\mathcal{V}_{W}}Y(V)$ (which lie in $W\times\mathbb{F}_{p}\cong\mathbb{F}_{p}^{k+1}$). Recall #$\#\mathcal{V}_{W}\gtrapprox p^{t_{2}}$ and #$\#Y(V)\sim p^{s}$. We see that $E_{W}$ is a $(s,(t_{2}-\varepsilon)_{+};k+1,k)$-set.
By Lemma 36, there exist dyadic numbers $p^{s_{1}(W)},p^{s_{2}(W)}$, so that $E_{W}$ contains a subset of form
@@
Here, $\Xi_{W}\subset W(\subset\mathbb{F}_{p}^{n-1}\times\mathbf{0})$ with #$\#\Xi_{W}\sim p^{s_{1}(W)}$; $\rho_{\xi}:=\xi+(\mathbf{0}^{n-1}\times\mathbb{F}_{p})$ and $Y_{W}(\rho_{\xi})\subset\rho_{\xi}$ with #$\#Y_{W}(\rho_{\xi})\sim p^{s_{2}(W)}$. We also have
@@
By pigeonholing on $(s_{1}(W),s_{2}(W))$ again, we can throw away some $W\in\mathcal{W}$ so that the cardinality of $\mathcal{W}$ is still $\gtrapprox p^{t_{1}}$, and assume there exist $s_{1},s_{2}$ such that
[…]
@@
(59) $\begin{split}E=\bigcup_{W\in\mathcal{W}}E_{W}&\supset\bigcup_{W\in\mathcal{W}}\bigcup_{\xi\in\Xi_{W}}Y_{W}(\rho_{\xi}).\end{split}$
Since #$\#Y_{W}(\rho_{\xi})\sim p^{s_{2}}$, we see that
#$\#E\gtrsim\#(\bigcup_{W\in\mathcal{W}}\Xi_{W})\cdot p^{s_{2}}.$
@@
#$\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{1},t_{1};n-1,k)-\varepsilon}p^{s_{2}}.$
The next lemma is a recursive inequality of $\mathbf{F}(s,t;n,k)$, which can be used to bound the right hand side of (57).
Lemma 37.
Suppose $(s,t;n,k)$ is admissible and $n\geq k+2$. Suppose there exist numbers $t_{1},t_{2},s_{1}$, such that
(60) $\begin{cases}t_{1}\in[0,(k+1)(n-k-1)];\\ t_{2}\in[0,k+1];\\ s_{1}\in[s,k];\end{cases}$
@@
(62) $\mathbf{F}(s_{1},t_{1};n-1,k)+\max\{\mathbf{F}(s,t_{2};k+1,k)-s_{1},0\}\geq\mathbf{F}(s,t;n,k).$
Let us quickly see how Lemma 35 and Lemma 37 combined to prove Proposition 34.
Proof of Proposition 34 assuming Lemma 35 and Lemma 37.
If $s=0$ or $t=0$, by Proposition 27 or 28, we are done. So, we suppose $s,t>0$.
@@
(63) #$\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;n,k)-C\varepsilon}$
@@
Suppose (53) holds. If $E$ is a $(s,t;n,k)$-set, by Lemma 35, there exist numbers $t_{1},t_{2},s_{1},s_{2}$, such that
(64) $\begin{cases}t_{1}\in[0,(k+1)(n-k-1)],t_{2}\in[0,k+1];\\ s_{1}\in[s-\varepsilon,k],s_{2}\in[0,1];\end{cases}$
@@
(65) $\begin{cases}t_{1}+t_{2}\geq t-\varepsilon;\\ s_{1}+s_{2}\geq\mathbf{F}(s,t_{2};k+1,k)-\varepsilon;\end{cases}$
@@
(66) #$\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{1},t_{1};n-1,k)+s_{2}-\varepsilon}\geq p^{\mathbf{F}(s_{1},t_{1};n-1,k)+\max\{\mathbf{F}(s,t_{2};k+1,k)-s_{1},0\}-2\varepsilon}.$
Let $t^{\prime}=t-\varepsilon$ which is $>0$ by assumption. Choose $0\leq t_{i}^{\prime}\leq t_{i}$ for $i=1,2$, so that $t_{1}^{\prime}+t_{2}^{\prime}=t^{\prime}$. Choose $s_{1}\leq s_{1}^{\prime}\leq s_{1}+\varepsilon$, so that $s_{1}^{\prime}\in[s,k]$. Then $t^{\prime},t_{1}^{\prime},t_{2}^{\prime},s_{1}^{\prime}$ satisfy the condition in Lemma 37, we obtain
@@
By monotonicity of $t\mapsto\mathbf{F}(s,t;n,k)$, Lemma 20 and Lemma 21, we have
#$\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;n,k)-C\varepsilon}.$
It remains to prove Lemma 37.
Proof of Lemma 37.
When $s=0$, we have $\mathbf{F}(0,t;n,k)=\max\{0,t-k(n-k)\}$. By Lemma 19, we have
@@
(67) $\mathbf{F}\geq\mathbf{F}(s,t_{2};k+1,k)+\mathbf{F}(s_{1},t_{1};n-1,k)-s_{1}.$
@@
(69) $\begin{cases}s=d+\sigma\\ t=(k-d-1)(n-k)+(d+2)m+\tau,\end{cases}$
@@
Here, the last line is the canonical expression for $t_{1}$ (see (9)), in which $20\leq m_{1}\leq n-k-$ 2, $\tau_{1}\in(0,d+\Delta+2]$.
@@
Proof of Theorem 7.
Finally, we see that Theorem 7 can be obtained from Proposition 29 and Proposition 34 using induction on $(n,k)$. ∎
6. Exceptional set estimate: lower bound
We prove Theorem 9 in this section. The result in this section has been studied by the author [4] under the Euclidean setting, the statement there is slightly different from this paper. We still provide the details for the finite field setting, though they are essentially the same.
Proof of Theorem 9.
Recall that $(n,k)=(2,1)$ and $\frac{a}{2}<s\leq\min\{1,a\}$ was considered in Proposition 16. We denote $\mathbf{A}_{s,a}\subset\mathbb{F}_{p}^{2}$ to be the set constructed there which satisfies #$\#\mathbf{A}_{s,a}\gtrsim p^{a}$, and
@@ 6. Exceptional set estimate: lower bound
Next, we prove for general $(n,k)$. Note that Type 1 and Type 4 are easy, so we consider Type 2 and Type 3. Recall the canonical expression as in Definition 14: $a=m+\beta,s=l+\gamma$.
Type 2 We have $l+1\leq m\leq n+l-k$. For $\gamma$, we only use $\gamma>\beta$. We first show the existence of $A$ with #$\#A\sim p^{a}$ such that #$\#E_{s}(A;n,k)\gtrsim p^{k(n-k)-(m-l)(k-l)}$.
Choose $A=\mathbb{F}_{p}^{m}\times I\times\mathbf{0}^{n-m-1}$, where $I\subset\mathbb{F}_{p}$ has cardinality $p^{\beta}$. For simplicity, we just write $A=\mathbb{F}_{p}^{m}\times I$. We want to find those $V\in G(n-k,\mathbb{F}_{p}^{n})$ such that #$\#(\pi_{V}^{*}(A))<p^{l+\gamma}$. We note that if $V$ satisfies #$\#\pi_{V}^{*}(\mathbb{F}_{p}^{m})\leq p^{l}$, then #$\#\pi_{V}^{*}(\mathbb{F}_{p}^{m}\times I)\leq p^{l+\beta}<p^{l+\gamma}$, which means $V\in E_{s}(A;n,k)$. This shows that
@@
By Lemma 18, the right hand side has cardinality $\sim p^{k(n-k)-(k-l)(m-l)}$.
Type 3, $\gamma\in(\frac{\beta}{2},\beta]$ We have $n-m-1\geq k-l,\gamma>\frac{\beta}{2}$. We show the existence of $A$ with #$\#A\sim p^{a}$ such that #$\#E_{s}(A;n,k)\gtrsim p^{k(n-k)-(m+1-l)(k-l)+2\gamma-\beta}$.
We choose $A=\mathbf{A}_{\gamma,\beta}\times\mathbb{F}_{p}^{m}\times\mathbf{0}^{n-m-2}$, where $\mathbf{A}_{\gamma,\beta}\subset\mathbb{F}_{p}^{2}$ is such that #$\#\mathbf{A}_{\gamma,\beta}\sim p^{\beta}$ and #$\#E_{\gamma}(\mathbf{A}_{\gamma,\beta};2,1)\gtrsim p^{2\gamma-\beta}$. We will use $\mathbb{F}_{p}^{2}$ to denote $\mathbb{F}_{p}^{2}\times\mathbf{0}^{n-2}$. We use $\mathbb{F}_{p}^{m}$ to denote $\mathbf{0}^{2}\times\mathbb{F}_{p}^{m}\times\mathbf{0}^{n-m-2}$. We use $\theta$ to denote the lines in $G(1,\mathbb{F}_{p}^{2})$. Define
@@
Let $\mathbf{E}_{\gamma,\beta}:=E_{\gamma}(\mathbf{A}_{\gamma,\beta};2,1)=\{\theta\in G(1,\mathbb{F}_{p}^{2}):\#(\pi_{\theta}^{*}(\mathbf{A}_{\gamma,\beta}))<p^{\gamma}\}$. We claim that if $\theta\in\mathbf{E}_{\gamma,\beta}$, then $\mathcal{V}_{\theta}\subset E_{l+\gamma}(A)$. In other words, $V\in\mathcal{V}_{\theta}$ satisfies #$\#(\pi_{V}^{*}(A))<p^{l+\gamma}$. We first note that by definition $\mathbf{A}_{\gamma,\beta}\subset\bigcup_{L\in\pi_{\theta}^{*}(\mathbf{A}_{\gamma,\beta})}L$, so
@@
#$\#(\pi_{V}^{*}(A))\leq\#(\pi_{\theta}^{*}(\mathbf{A}_{\gamma,\beta}))\cdot\#(\pi_{V}^{*}(\theta\oplus\mathbb{F}_{p}^{m}))<p^{\gamma+l}.$
@@
To make the argument precise, we first apply Lemma 18 to see
#$\#\mathcal{V}_{\theta}\sim p^{k(n-k)-(m+1-l)(k-l)}.$
To calculate #$\#\mathcal{W}_{\theta}$, we apply Lemma 18 again to see
#$\#\{V\in G(n-k,n):\dim(\pi_{V}^{*}(\mathbb{F}_{p}^{m}))\leq p^{l-1}\}\sim p^{k(n-k)-(m-l+1)(k-l+1)},$
@@
#$\#\{V\in G(n-k,n):\dim(\pi_{V}^{*}(\mathbb{F}_{p}^{2}\times\mathbb{F}_{p}^{m}))\leq p^{l}\}\sim p^{k(n-k)-(m+2-l)(k-l)}.$
Therefore, #$\#\mathcal{W}_{\theta}\ll\#\mathcal{V}_{\theta}$ (when $p$ is large enough) and hence
#$\#\mathcal{U}_{\theta}\sim\#\mathcal{V}_{\theta}\sim p^{k(n-k)-(m+1-l)(k-l)}.$
@@
#$\#(\pi_{V}^{*}(\theta\oplus\mathbb{F}_{p}^{m}))\leq p^{l},$
#$\#(\pi_{V}^{*}(\theta^{\prime}\oplus\mathbb{F}_{p}^{m}))\leq p^{l},$
#$\#(\pi_{V}^{*}(\mathbb{F}_{p}^{m}))=p^{l},$
#$\#(\pi_{V}^{*}(\mathbb{F}_{p}^{2}\times\mathbb{F}_{p}^{m}))=p^{l+1}.$
@@
which has cardinality $=p^{l}$, contradicting the fourth condition.
We have shown that $E_{l+\gamma}(A;n,k)\supset\bigsqcup_{\theta\in\mathbf{E}_{\gamma,\beta}}\mathcal{U}_{\theta}$, which has cardinality #$\#\mathbf{E}_{\gamma,\beta}\cdot\#\mathcal{U}_{\theta}$
@@
Type 2, $\gamma\in(\frac{\beta+1}{2},1]$ The trick is to write $a=(m-1)+(\beta+1)=:m^{\prime}+\beta^{\prime}$. Note that $l\leq m^{\prime}\leq n+l-k-1$ and $\gamma>\frac{\beta^{\prime}}{2}$. We apply the construction in “Type 3, $\gamma\in(\frac{\beta}{2},\beta]$”, where we only used $\gamma>\frac{\beta}{2}$. We obtain a $A$ with #$\#A\sim p^{a}$ such that
#$\#E_{s}(A;n,k)\gtrsim p^{k(n-k)-(m^{\prime}+1-l)(k-l)+2\gamma-\beta^{\prime}}=p^{k(n-k)-(m-l)(k-l)+2\gamma-(\beta+1)}.$
Type 3, $\gamma\in(0,\frac{\beta}{2}]$ The trick is to make $a$ bigger. We write $a=m+\beta\leq(m+1)+0=:m^{\prime}+\beta^{\prime}$. One can check $l+1\leq m^{\prime}\leq n+l-k$ and $\gamma>\beta^{\prime}$. We apply the construction in Type 2 with $(m^{\prime},\beta^{\prime})$. We find $A^{\prime}$ with #$\#A^{\prime}\sim p^{m+1}$ such that
#$\#E_{s}(A^{\prime};n,k)\gtrsim p^{k(n-k)-(m+1-l)(k-l)}.$
Choosing any subset $A\subset A^{\prime}$ with #$\#A\sim p^{a}$, we get
#$\#E_{s}(A;n,k)\geq\#E_{s}(A^{\prime};n,k)\gtrsim p^{k(n-k)-(m+1-l)(k-l)}.$
7. Exceptional set estimate: Upper bound I

### v2

@@ 5 Furstenberg set problem: Lower bound II
Return to $E_{W}:=\bigcup_{V\in\mathcal{V}_{W}}Y(V)$ (which lies in $W\times\mathbb{F}_{p}\cong\mathbb{F}_{p}^{k+1}$). Recall $\mathcal{V}_{W}\gtrapprox p^{t_{2}}$ and $Y(V)\sim p^{s}$. We see that $E_{W}$ is a $(s,(t_{2}-\varepsilon)_{+};k+1,k)$-set.
By Lemma 5.3, there exist dyadic numbers $p^{s_{1}(W)},p^{s_{2}(W)}$, so that $E_{W}$ contains a subset of form
@@
Here, $\Xi_{W}\subset W(\subset\mathbb{F}_{p}^{n-1}\times\mathbf{0})$ with $\Xi_{W}\sim p^{s_{1}(W)}$; $\rho_{\xi}:=\xi+(\mathbf{0}^{n-1}\times\mathbb{F}_{p})$ and $Y_{W}(\rho_{\xi})\subset\rho_{\xi}$ with $Y_{W}(\rho_{\xi})\sim p^{s_{2}(W)}$. We also have
@@
By pigeonholing on $(s_{1}(W),s_{2}(W))$ again, we can throw away some $W\in\mathcal{W}$ so that the cardinality of $\mathcal{W}$ is still $\gtrapprox p^{t_{1}}$, and assume that there exist $s_{1},s_{2}$ such that
$(s_{1}(W),s_{2}(W))=(s_{1},s_{2}),\ \textup{~for~all~}W\in\mathcal{W},$
@@
$\begin{split}E=\bigcup_{W\in\mathcal{W}}E_{W}&\supset\bigcup_{W\in\mathcal{W}}\bigcup_{\xi\in\Xi_{W}}Y_{W}(\rho_{\xi}).\end{split}$ (59)
Since $Y_{W}(\rho_{\xi})\sim p^{s_{2}}$, we see that
$E\gtrsim\#(\bigcup_{W\in\mathcal{W}}\Xi_{W})\cdot p^{s_{2}}.$
@@
$E\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{1},t_{1};n-1,k)-\varepsilon}p^{s_{2}}.$
The next lemma is a recursive inequality of $\mathbf{F}(s,t;n,k)$, which can be used to bound the right-hand side of (57).
Lemma 5.4.
Suppose $(s,t;n,k)$ is admissible and $n\geq k+2$. Suppose that there exist numbers $t_{1},t_{2},s_{1}$, such that
$\begin{cases}t_{1}\in[0,(k+1)(n-k-1)];\\ t_{2}\in[0,k+1];\\ s_{1}\in[s,k];\end{cases}$ (60)
@@
$\mathbf{F}(s_{1},t_{1};n-1,k)+\max\{\mathbf{F}(s,t_{2};k+1,k)-s_{1},0\}\geq\mathbf{F}(s,t;n,k).$ (62)
Let us quickly see how Lemma 5.2 and Lemma 5.4 combined to prove Proposition 5.1.
Proof of Proposition 5.1 assuming Lemma 5.2 and Lemma 5.4.
If $s=0$ or $t=0$, by Proposition 4.1 or 4.2, we are done. So, we suppose $s,t>0$.
@@
$E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;n,k)-C\varepsilon}$ (63)
@@
Suppose (53) holds. If $E$ is a $(s,t;n,k)$-set, by Lemma 5.2, there exist numbers $t_{1},t_{2},s_{1},s_{2}$, such that
$\begin{cases}t_{1}\in[0,(k+1)(n-k-1)],t_{2}\in[0,k+1];\\ s_{1}\in[s-\varepsilon,k],s_{2}\in[0,1];\end{cases}$ (64)
@@
$\begin{cases}t_{1}+t_{2}\geq t-\varepsilon;\\ s_{1}+s_{2}\geq\mathbf{F}(s,t_{2};k+1,k)-\varepsilon;\end{cases}$ (65)
@@
$E\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{1},t_{1};n-1,k)+s_{2}-\varepsilon}\geq p^{\mathbf{F}(s_{1},t_{1};n-1,k)+\max\{\mathbf{F}(s,t_{2};k+1,k)-s_{1},0\}-2\varepsilon}.$ (66)
Let $t^{\prime}=t-\varepsilon$ which is >0>0 by assumption. Choose $0\leq t_{i}^{\prime}\leq t_{i}$ for $i=1,2$, so that $t_{1}^{\prime}+t_{2}^{\prime}=t^{\prime}$. Choose $s_{1}\leq s_{1}^{\prime}\leq s_{1}+\varepsilon$, so that $s_{1}^{\prime}\in[s,k]$. Then $t^{\prime},t_{1}^{\prime},t_{2}^{\prime},s_{1}^{\prime}$ satisfy the condition in Lemma 5.4, we obtain
@@
By the monotonicity of $t\mapsto\mathbf{F}(s,t;n,k)$, Lemma 2.11 and Lemma 2.12, we have
$E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;n,k)-C\varepsilon}.$
It remains to prove Lemma 5.4.
Proof of Lemma 5.4.
When $s=0$, we have $\mathbf{F}(0,t;n,k)=\max\{0,t-k(n-k)\}$. By Lemma 2.10, we have
@@
$\mathbf{F}\geq\mathbf{F}(s,t_{2};k+1,k)+\mathbf{F}(s_{1},t_{1};n-1,k)-s_{1}.$ (67)
@@
$\begin{cases}s=d+\sigma\\ t=(k-d-1)(n-k)+(d+2)m+\tau,\end{cases}$ (69)
@@
Here, the last line is the canonical expression for $t_{1}$ (see (9)), in which $0\leq m_{1}\leq n-k-2$, $\tau_{1}\in(0,d+\Delta+2]$.
@@
Proof of Theorem 1.8.
Finally, we see that Theorem 1.8 can be obtained from Proposition 4.3 and Proposition 5.1 using induction on $(n,k)$. ∎
6 Exceptional set estimate: lower bound
We prove Theorem 1.10 in this section. The result in this section has been studied by the author [8] under the Euclidean setting, and the statement there is slightly different from this paper. We still provide the details for the finite-field setting, though they are essentially the same.
Proof of Theorem 1.10.
Recall that $(n,k)=(2,1)$ and $\frac{a}{2}<s\leq\min\{1,a\}$ was considered in Proposition 2.6. We denote $\mathbf{A}_{s,a}\subset\mathbb{F}_{p}^{2}$ to be the set constructed there which satisfies $\mathbf{A}_{s,a}\gtrsim p^{a}$, and
@@ 6 Exceptional set estimate: lower bound
Next, we prove for general $(n,k)$. Note that Type 1 and Type 4 are easy, so we consider Type 2 and Type 3. Recall the canonical expression as in Definition 2.4: $a=m+\beta,s=l+\gamma$.
Type 2We have $l+1\leq m\leq n+l-k$. For $\gamma$, we only use $\gamma>\beta$. We first show the existence of $A$ with $A\sim p^{a}$ such that $E_{s}(A;n,k)\gtrsim p^{k(n-k)-(m-l)(k-l)}$.
Choose $A=\mathbb{F}_{p}^{m}\times I\times\mathbf{0}^{n-m-1}$, where $I\subset\mathbb{F}_{p}$ has cardinality $p^{\beta}$. For simplicity, we just write $A=\mathbb{F}_{p}^{m}\times I$. We want to find those $V\in G(n-k,\mathbb{F}_{p}^{n})$ such that $(\pi_{V}^{*}(A))<p^{l+\gamma}$. We note that if $V$ satisfies $\pi_{V}^{*}(\mathbb{F}_{p}^{m})\leq p^{l}$, then $\pi_{V}^{*}(\mathbb{F}_{p}^{m}\times I)\leq p^{l+\beta}<p^{l+\gamma}$, which means $V\in E_{s}(A;n,k)$. This shows that
@@
By Lemma 2.8, the right hand side has cardinality $\sim p^{k(n-k)-(k-l)(m-l)}$.
Type 3, $\gamma\in(\frac{\beta}{2},\beta]$We have $n-m-1\geq k-l,\gamma>\frac{\beta}{2}$. We show the existence of $A$ with $A\sim p^{a}$ such that $E_{s}(A;n,k)\gtrsim p^{k(n-k)-(m+1-l)(k-l)+2\gamma-\beta}$.
We choose $A=\mathbf{A}_{\gamma,\beta}\times\mathbb{F}_{p}^{m}\times\mathbf{0}^{n-m-2}$, where $\mathbf{A}_{\gamma,\beta}\subset\mathbb{F}_{p}^{2}$ is such that $\mathbf{A}_{\gamma,\beta}\sim p^{\beta}$ and $E_{\gamma}(\mathbf{A}_{\gamma,\beta};2,1)\gtrsim p^{2\gamma-\beta}$. We will use $\mathbb{F}_{p}^{2}$ to denote $\mathbb{F}_{p}^{2}\times\mathbf{0}^{n-2}$. We use $\mathbb{F}_{p}^{m}$ to denote $\mathbf{0}^{2}\times\mathbb{F}_{p}^{m}\times\mathbf{0}^{n-m-2}$. We use $\theta$ to denote the lines in $G(1,\mathbb{F}_{p}^{2})$. Define
@@
Let $\mathbf{E}_{\gamma,\beta}:=E_{\gamma}(\mathbf{A}_{\gamma,\beta};2,1)=\{\theta\in G(1,\mathbb{F}_{p}^{2}):\#(\pi_{\theta}^{*}(\mathbf{A}_{\gamma,\beta}))<p^{\gamma}\}$. We claim that if $\theta\in\mathbf{E}_{\gamma,\beta}$, then $\mathcal{V}_{\theta}\subset E_{l+\gamma}(A)$. In other words, $V\in\mathcal{V}_{\theta}$ satisfies $(\pi_{V}^{*}(A))<p^{l+\gamma}$. We first note that by definition $\mathbf{A}_{\gamma,\beta}\subset\bigcup_{L\in\pi_{\theta}^{*}(\mathbf{A}_{\gamma,\beta})}L$, so
@@
$(\pi_{V}^{*}(A))\leq\#(\pi_{\theta}^{*}(\mathbf{A}_{\gamma,\beta}))\cdot\#(\pi_{V}^{*}(\theta\oplus\mathbb{F}_{p}^{m}))<p^{\gamma+l}.$
@@
To make the argument precise, we first apply Lemma 2.8 to see
$\mathcal{V}_{\theta}\sim p^{k(n-k)-(m+1-l)(k-l)}.$
To calculate $\mathcal{W}_{\theta}$, we apply Lemma 2.8 again to see
$\{V\in G(n-k,n):\dim(\pi_{V}^{*}(\mathbb{F}_{p}^{m}))\leq p^{l-1}\}\sim p^{k(n-k)-(m-l+1)(k-l+1)},$
@@
$\{V\in G(n-k,n):\dim(\pi_{V}^{*}(\mathbb{F}_{p}^{2}\times\mathbb{F}_{p}^{m}))\leq p^{l}\}\sim p^{k(n-k)-(m+2-l)(k-l)}.$
Therefore, $\mathcal{W}_{\theta}\ll\#\mathcal{V}_{\theta}$ (when $p$ is large enough) and hence
$\mathcal{U}_{\theta}\sim\#\mathcal{V}_{\theta}\sim p^{k(n-k)-(m+1-l)(k-l)}.$
@@
$\#(\pi_{V}^{*}(\theta\oplus\mathbb{F}_{p}^{m}))\leq p^{l},$
$\#(\pi_{V}^{*}(\theta^{\prime}\oplus\mathbb{F}_{p}^{m}))\leq p^{l},$
$\#(\pi_{V}^{*}(\mathbb{F}_{p}^{m}))=p^{l},$
$\#(\pi_{V}^{*}(\mathbb{F}_{p}^{2}\times\mathbb{F}_{p}^{m}))=p^{l+1}.$
@@
which has cardinality $p^{l}$, contradicting the fourth condition.
We have shown that $E_{l+\gamma}(A;n,k)\supset\bigsqcup_{\theta\in\mathbf{E}_{\gamma,\beta}}\mathcal{U}_{\theta}$, which has cardinality $\mathbf{E}_{\gamma,\beta}\cdot\#\mathcal{U}_{\theta}$
@@
Type 2, $\gamma\in(\frac{\beta+1}{2},1]$The trick is to write $a=(m-1)+(\beta+1)=:m^{\prime}+\beta^{\prime}$. Note that $l\leq m^{\prime}\leq n+l-k-1$ and $\gamma>\frac{\beta^{\prime}}{2}$. We apply the construction in “Type 3, $\gamma\in(\frac{\beta}{2},\beta]$”, where we only used $\gamma>\frac{\beta}{2}$. We obtain a $A$ with $A\sim p^{a}$ such that
$E_{s}(A;n,k)\gtrsim p^{k(n-k)-(m^{\prime}+1-l)(k-l)+2\gamma-\beta^{\prime}}=p^{k(n-k)-(m-l)(k-l)+2\gamma-(\beta+1)}.$
Type 3, $\gamma\in(0,\frac{\beta}{2}]$The trick is to make $a$ bigger. We write $a=m+\beta\leq(m+1)+0=:m^{\prime}+\beta^{\prime}$. One can check $l+1\leq m^{\prime}\leq n+l-k$ and $\gamma>\beta^{\prime}$. We apply the construction in Type 2 with $(m^{\prime},\beta^{\prime})$. We find $A^{\prime}$ with $A^{\prime}\sim p^{m+1}$ such that
$E_{s}(A^{\prime};n,k)\gtrsim p^{k(n-k)-(m+1-l)(k-l)}.$
Choosing any subset $A\subset A^{\prime}$ with $A\sim p^{a}$, we get
$E_{s}(A;n,k)\geq\#E_{s}(A^{\prime};n,k)\gtrsim p^{k(n-k)-(m+1-l)(k-l)}.$
7 Exceptional set estimate: Upper bound I